OpenStudy (anonymous):

someone solve this proportion please 5/13=v/9

OpenStudy (anonymous):

5/13=v/9 --> 13v=45 --> v=3 6/13 (as in 45/13)

OpenStudy (lalaly):

\[\frac{5}{13} = \frac{v}{9}\]\[5 \times 9 = 13 \times v\]\[45=13v\]\[v=\frac{45}{13}\]

OpenStudy (anonymous):

could you round the answer to the neares tenth?? plz

OpenStudy (anonymous):

3.5 i think:)

OpenStudy (lalaly):

yeh thts right